# Can Tool-augmented Large Language Models be Aware of Incomplete Conditions?

### v2

@@ Abstract
Recent advancements in integrating large language models (LLMs) with tools have allowed the models to interact with real-world environments. However, these tool-augmented LLMs often encounter incomplete scenarios when users provide partial information or the necessary tools are unavailable. Recognizing and managing such scenarios is crucial for LLMs to ensure their reliability, but this exploration remains understudied. This study examines whether LLMs can identify incomplete conditions and appropriately determine when to refrain from using tools. To this end, we address a dataset by manipulating instances from two datasets by removing necessary tools or essential information for tool invocation. We confirm that most LLMs are challenged to identify the additional information required to utilize specific tools and the absence of appropriate tools. We further analyze model behaviors in different environments and compare their performance against humans. Our research can contribute to advancing reliable LLMs by addressing scenarios that commonly arise during interactions between humans and LLMs. [^1] Our code and dataset will be made publicly available.
@@ 1 Introduction
Recently, there has been significant improvement in integrating large language models (LLMs) with tools (li-etal-2023-api; qin2023toolllm; patil2023gorilla; schick2024toolformer; lu2024chameleon; hao2024toolkengpt). These tool-augmented LLMs can perceive up-to-date information, acquire real-world interaction capabilities, and perform complex tasks (tool-anyway). The tool-augmented models have enhanced user experiences across various applications (autogpt; hong2023metagpt).
@@
LLMs should select and use the necessary tools when they interact with users. However, users often lack comprehensive knowledge about all available tools, making it challenging to initially provide the required information to models. Furthermore, the necessary tools may not be provided to models or even do not exist to address user requests. When LLM agents use various tools, it is common to encounter issues where the necessary tools are absent, or the user’s input is incomplete, lacking the information needed to use them. Despite significant advances in tool-augmented LLM, research has been conducted on the premise of only a complete scenario (huang2023metatool).
Our work investigates whether tool-augmented LLMs can recognize incomplete conditions where no appropriate tool is available for the user’s request, or the given information is insufficient to utilize the tool, as depicted in Fig. 1. To this end, we construct a dataset by manipulating instances from datasets for tool-augmented LLMs (qin2023toolllm; li-etal-2023-api) to simulate the scenarios. Further human verification ensures only instances where tools could not be used are included.
Based on our constructed dataset, we evaluate whether LLMs can recognize conditions where desired tool invocation is not possible. Our experimental results show that most LLMs struggle to identify the additional information required to use specific tools or to recognize when appropriate tools are unavailable. Specifically, LLMs perform worse when users provide incomplete information or when real-world tools are involved. To further understand these challenges, we conduct comprehensive analyses, including human evaluation and assessing the impact of providing API call feedback to models. We also explore the predictive distributions of models and their ability to explain decisions in incomplete scenarios. We believe our findings contribute to advancing the reliability of tool-augmented LLMs in practical applications. Our contributions can be summarized as follows.
  • We construct a dataset by manipulating tool-use datasets to simulate incomplete conditions in real-world applications.
  • We evaluate the ability of LLMs to recognize when desired tool invocation is not possible, showing that LLMs struggle with identifying the need for additional information or recognizing the absence of appropriate tools.
  • We conduct comprehensive analyses, including human evaluations and assessments of API call feedback, to explore model behavior in incomplete scenarios, providing insights that contribute to improving the reliability of tool-augmented LLMs.
@@ 2 Related Work
Recent work has focused on augmenting LLMs with external tools to enhance their capabilities (tool-anyway). These tools range from simple retrieval systems (chen-etal-2017-reading) and arithmetic operations (schick2024toolformer) to more complex programming languages (llm-toolmaker) and real-world APIs (xu2023tool; yuan2024easytool; guo2024stabletoolbench). Various strategies and benchmarks have been developed to improve how effectively LLMs use these tools. For example, huang2023metatool evaluate how well LLMs decide when and which tools to use, and ye2024rotbench assess their robustness when noisy tool descriptions are present. huang-etal-2024-planning-creation investigate LLMs’ use of tools in real-world scenarios, and ye-etal-2024-toolsword address safety challenges in tool learning. Although berkeley-function-calling-leaderboard explored how LLMs perform in the absence of necessary tools, their dataset is relatively small compared to ours and considers only a single API candidate, which does not adequately reflect the complexities of the real world. Additionally, our work extends incomplete scenarios by considering where tool invocation is impossible due to missing or insufficient information. While research in question answering and retrieval systems has addressed scenarios involving irrelevant external knowledge or ambiguous user requests (feng-etal-2020-none; min-etal-2020-ambigqa; kamath-etal-2020-selective; cole-etal-2023-selectively; adaptive-rag), the challenge of LLMs recognizing when tool usage is impractical has been relatively underexplored. Our study advances this area by investigating how LLMs recognize situations where tools are unavailable or cannot be correctly invoked due to incomplete information.
@@ 3 How to Evaluate the Awareness of Tool-augmented LLMs
To simulate incomplete scenarios where necessary tools are unavailable or users provide partial information, we manipulate instances from a test set of two benchmarks: APIBank (li-etal-2023-api) and ToolBench (qin2023toolllm). These datasets are designed to evaluate how effectively LLMs can respond to user requests using APIs. [^2] We refer to Tool and API interchangeably. In their original form, each instance in these datasets has an available API that can address the user’s request, with sufficient information provided to invoke the API. This section first provides an overview of the source datasets (section 3.1) and then introduces our automated data manipulation strategy to simulate incomplete scenarios for tool usage (section 3.2). Finally, we describe the extensive human verification process that ensures the validity of the manipulated dataset (section 3.3). Fig. 2 illustrates the overall dataset construction pipeline.
@@ 3.1 Data Source
We construct a new dataset by leveraging two existing datasets: APIBank (li-etal-2023-api) and ToolBench (qin2023toolllm). APIBank consists of 73 APIs and 314 manually annotated multi-turn conversations between humans and AI systems. We select 450 instances from the test split, excluding those that require a tool-retrieval module. ToolBench is based on 16,000 real-world APIs across 49 categories from RapidAPI Hub. [^3] https://rapidapi.com/hub We utilize 764 test instances in different categories (single-tool instructions (I1), intra-category multi-tool instructions (I2), and intra-collection multi-tool instructions (I3)), as filtered by guo2024stabletoolbench. These instances are used as the original data for manipulation.
@@ 3.2 Simulating Incomplete Scenarios with Instance Manipulation
We simulate two types of incomplete scenarios in which LLMs cannot properly invoke tools by manipulating the original dataset instances: (1) replacing relevant APIs with irrelevant ones, and (2) partially removing user utterances.
Figure 2: Dataset construction pipeline.
@@
We use a dense retriever (gao-etal-2021-simcse) to replace the appropriate APIs that can handle the user’s requests with similar but irrelevant APIs. This manipulation ensures that the desired tool is unavailable, simulating a scenario where LLMs must decide not to use any of the provided tools. Specifically, the relevant APIs in the original instance are replaced with other APIs from separate API pools. To find semantically similar APIs, we use the sentence encoder proposed by gao-etal-2021-simcse.444princeton-nlp/sup-simcse-roberta-large We concatenate the name and description of each API, convert this text into a fixed-size vector, and then select one of the most similar APIs by calculating cosine similarity between the relevant API and all available APIs in the dataset. Note that the API pools for this retrieval process are constructed separately for each dataset. Examples of successful and failed instances of this manipulation are shown in Fig. LABEL:fig:app:apibank_api_replacement_success to Fig. LABEL:fig:app:toolbench_api_replacement_wrong.
@@
We partially remove user utterances in a conversation to mimic scenarios where (1) the request is unclear, making it impossible to use APIs, or (2) the request is clear but lacks necessary information for API invocation. This manipulation aims to remove essential information from user utterances, rendering appropriate API invocation infeasible. We automate this process using a proprietary LLM (i.e., GPT-4 (achiam2023gpt)) to generate naturally corrupted dialogues. The model is instructed to identify and remove critical information required for tool invocation, and we enhance the quality of this process with reasoning prompts and five manually designed few-shot samples. Examples of successful and failed instances are shown in Fig. LABEL:fig:app:apibank_utterance_removal_success to Fig. LABEL:fig:app:toolbench_utterance_removal_wrong.
@@ 3.3 Data Verification
To ensure the validity of our dataset, we manually reviewed all instances to remove cases where user requests could still be handled by the provided APIs or where the manipulations resulted in unnatural conversations. This manual verification was necessary because our preliminary analysis revealed that automatic generation sometimes failed to create truly incomplete scenarios. The verification process was conducted by two authors with relevant academic backgrounds, who actively discussed and refined the filtering criteria. This rigorous process ensured that the final dataset includes only instances that genuinely represent incomplete scenarios, providing a robust foundation for evaluating tool-augmented LLMs.
Specifically, we applied a detailed validation process to ensure that only incomplete conditions remained in the manipulated instances. The criteria for verifying Relevant API Replacement were: (1) the user’s request could still be fulfilled using alternative APIs, and (2) the request could be handled using non-replaced APIs, particularly in ToolBench instances. For Partial Removal of User Utterance, the criteria included: (1) only non-essential information was removed, allowing the APIs to fulfill the user’s request, (2) one of the user’s multiple requests was removed, but the remaining requests could still be resolved through existing API invocations, (3) the conversation context became unnatural or grammatically incorrect, (4) the result was identical to the original instance, and (5) it became unnatural for the model to execute an API call in the subsequent turn because the last utterance in the conversation no longer pertained to the user.
@@ 3.4 Dataset Statistics
Table 1: Dataset Statistics. The API Replacement and Utterance Removal denote Relevant API Replacement and Partial Removal of User Utterance, respectively. The Avg. Turns denote the number of utterances in a conversation. Avg. APIs denote the number of APIs in the instance. Avg. API Length denotes the string length of each API description.
@@ 4 Experimental Setup
We set the task where LLMs recognize incomplete conditions as a binary classification problem. We ask the LLM to determine whether APIs can be invoked to fulfill user requests based on conversations and available APIs. To assess whether LLMs can recognize both complete and incomplete conditions, we utilize the instances of two datasets(APIBank and ToolBench) and corresponding manipulated samples. LLMs should answer "Yes" or "No" for the original and manipulated instances, respectively.
@@
For the experiments, we adopt both open-source and proprietary LLMs. The open-source models used are Llama-2 (13B-chat, 70B-chat) (touvron2023llama), Vicuna-v1.5 (13B) (zheng2024judging), and Mistral-Instruct-v0.2 (7B) (jiang2023mistral). The following proprietary LLMs are used: Claude-3-Haiku (anthropic2024claude), GPT-3.5-Turbo (openai2023chatgpt), and GPT-4 (achiam2023gpt).
Implementation Details
All open-source LLMs are implemented with the Transformers library (wolf-etal-2020-transformers). The quantized checkpoints of Llama-2-70B-hf and Llama-2-70B-chat are used for experiments (lin2023awq). For proprietary models, we use claude-3-haiku-20240307, gpt-3.5-turbo-0125, and gpt-4-0613 for Claude-3-Haiku, GPT-3.5-Turbo, and GPT-4. The temperature is set to 0 across all models. We use two original and two manipulated instances for the evaluation with few-shot examples. In experiments with instances from ToolBench, we select models with a context length exceeding 8192 tokens to handle the frequently encountered lengthy API descriptions.
Mistral-7B Claude-3 GPT-3.5-T GPT-4 API Replacement Acc. 64.69/67.91 68.42/61.90 66.03/64.40 79.99/82.29 F1 47.98/52.13 60.78/36.81 54.12/45.14 74.68/80.93 Utterance Removal Acc. 51.29/51.42 57.59/51.25 52.71/57.90 73.21/77.44 F1 14.12/12.33 40.05/5.41 19.80/28.64 58.20/72.30
Table 2: Performance evaluation results of LLM by manipulation type. The accuracy (Acc.) and F1 score (F1) are used for evaluation metrics. Both the zero-shot and four-shot performance are presented in a format of (0-shot/4-shot). The highest and the second-highest scores in each metric are highlighted in bold and underlined. The performance of models on APIBank and ToolBench are presented in Table 3 and Table 4, respectively.
Llama-2-13B Llama-2-70B Vicuna-13B Mistral-7B Claude-3 GPT-3.5-T GPT-4 API Replacement Acc. 53.75/57.07 59.66/58.30 61.87/66.17 69.25/74.91 68.27/65.19 67.77/70.85 87.58/85.49 F1 14.55/32.50 36.19/32.34 45.99/64.15 61.66/69.28 62.61/46.30 62.03/64.99 87.30/84.79 Utterance Removal Acc. 51.29/52.15 52.15/51.12 51.12/54.04 52.32/52.32 61.96/52.50 55.94/65.40 89.33/88.64 F1 5.98/18.71 16.77/12.88 18.39/44.95 22.84/20.17 52.27/9.80 36.63/55.23 89.27/88.70
Table 3: The performance of models on instances from APIBank.
Mistral-7B Claude-3 GPT-3.5-T GPT-4 API Replacement Acc. 60.13/60.90 68.57/58.60 64.29/57.94 72.40/79.08 F1 34.30/34.97 58.94/27.31 46.20/25.29 62.05/77.07 Utterance Removal Acc. 50.26/50.52 53.22/50.00 49.48/50.39 57.09/66.24 F1 5.39/4.48 27.83/1.02 2.97/2.04 27.13/55.89
Table 4: The performance of models on instances from ToolBench.
5 Results
Main Results
Table 2 presents the performance of different LLMs averaged on APIBank and ToolBench. LLMs usually struggle to identify incomplete conditions, and proprietary models perform better than Mistral-7B. In most cases, models are more vulnerable to Utterance Removal compared to API Replacement. For instance, in a zero-shot experiment, the F1 score for Mistral-7B decreases from 47.98% to 14.12%. Similarly, GPT-3.5-Turbo drops from 54.12% to 19.80%, and GPT-4 declines from 74.68% to 58.2%.
Impacts of Few-shot Examples
We examine the impact of providing the model with demonstration examples on its performance. To this end, we provide manually curated few-shot examples, ensuring the classes are evenly balanced. Surprisingly, there is no clear tendency indicating either performance improvement or degradation when using few-shot learning samples. We suspect that using the tool requires significant reasoning skills and involves various scenarios, making it difficult to improve the incomplete condition ability with only a few examples (khot2022decomposed).
Impacts of Data Sources
We next investigate performance by categorizing it based on data sources in Table 3 and Table 4. Among the open-source models, Mistral-7B performs the best on instances from APIBank. At the same time, the open-source model is particularly vulnerable compared to the proprietary model in scenarios where the user did not provide sufficient information to call APIs. The performance of models on instances from ToolBench is generally lower compared to their performance on APIBank. Even the most powerful model known to be, GPT-4, shows a low F1 score in a zero-shot experiment, with 62.05% in the API Replacement and 27.13% in the Utterance Removal. These tendencies can stem from the complexity of the API descriptions in ToolBench, which consists of real-world APIs that frequently contain unnecessary information or omit essential details (yuan2024easytool).

### v3

@@ Abstract
Recent advancements in integrating large language models (LLMs) with tools have allowed the models to interact with real-world environments. However, these tool-augmented LLMs often encounter incomplete scenarios when users provide partial information or the necessary tools are unavailable. Recognizing and managing such scenarios is crucial for LLMs to ensure their reliability, but this exploration remains understudied. This study examines whether LLMs can identify incomplete conditions and appropriately determine when to refrain from using tools. To this end, we address a dataset by manipulating instances from two datasets by removing necessary tools or essential information for tool invocation. Our experiments show that LLMs often struggle to identify the absence of information required to utilize specific tools and recognize the absence of appropriate tools. We further analyze model behaviors in different environments and compare their performance against humans. Our research can contribute to advancing reliable LLMs by addressing common scenarios during interactions between humans and LLMs. [^1] Our code and dataset will be publicly available.
@@ 1 Introduction
Recently, there has been significant improvement in integrating large language models (LLMs) with tools (Li et al., 2023; Qin et al., 2023; Patil et al., 2023; Schick et al., 2024; Hao et al., 2024). These tool-augmented LLMs can perceive up-to-date information, acquire real-world interaction capabilities, and perform complex tasks (Wang et al., 2024), enhancing user experiences across various applications (Yang et al., 2023; Hong et al., 2023).
@@
Despite their capabilities, tool-augmented LLMs often face scenarios where users lack comprehensive knowledge about available tools or the necessary tools are absent. When LLM agents interact with users, they frequently encounter situations with missing tools or incomplete information needed for tool usage. While significant advances have been made in tool-augmented LLMs, existing research has primarily focused on complete scenarios (Huang et al., 2023; Zhang et al., 2024b).
We investigate whether tool-augmented LLMs can recognize incomplete conditions where no appropriate tool is available or the provided information is insufficient for tool utilization, as depicted in Fig.1. We construct a dataset by manipulating instances from existing tool-augmented LLM datasets (Qin et al., 2023; Li et al., 2023), with human verification ensuring valid instances.
Our experimental results show varying performance patterns across different types of incomplete scenarios: while larger models generally demonstrate strong performance in recognizing tool unavailability, the performance gaps become more pronounced when handling insufficient information for tool usage, particularly in real-world applications. To better understand these capabilities, we conduct comprehensive analyses, including human evaluation and Chain-of-Thought (CoT) prompting (Wei et al., 2022), few-shot learning approaches, and API call feedback effects. Our findings show that while larger models approach or exceed human-level performance, the effectiveness of enhancement strategies varies across model sizes and tasks.
Our contributions are: (1) We construct a dataset simulating incomplete conditions by manipulating tool-use datasets. (2) We evaluate LLMs’ ability to recognize impossible tool invocations, revealing their struggles with identifying missing information. (3) We provide comprehensive analyses of model behavior in incomplete scenarios.
@@ 2 Related Work
Recent research has explored the capability enhancement of LLMs with external tools (Wang et al., 2024; Gu et al., 2024; Cai et al., 2024; Fan et al., 2024), ranging from basic retrieval systems (Chen et al., 2017) and arithmetic operations (Schick et al., 2024) to complex programming languages (Gou et al., 2024; Zhang et al., 2024a) and APIs (Xu et al., 2023; Yuan et al., 2024; Guo et al., 2024). Various benchmarks have been developed to evaluate LLMs’ tool usage capabilities, including tool selection timing (Huang et al., 2023), robustness to noisy descriptions (Ye et al., 2024b), error handling (Sun et al., 2024), and safety considerations (Ye et al., 2024a).
Although some studies have examined the awareness of LLMs when the necessary tools are not provided, the existing work has significant limitations. Previous studies focused on limited tool scenarios (Ning et al., 2024; Yan et al., 2024), assumed perfect tool utilization plan is given (Huang et al., 2024), or overlooked situations with similar but incorrect tools (Zhang et al., 2024b). Our work addresses these gaps by exploring more realistic scenarios, including situations where incorrect but similar tools are provided or where users fail to supply the necessary information for tool invocation. Although research in question answering has explored handling irrelevant knowledge and ambiguous requests (Feng et al., 2020; Min et al., 2020; Kamath et al., 2020; Cole et al., 2023; Jeong et al., 2024), LLMs’ ability to recognize impractical tool usage remains underexplored.
@@ 3 How to Evaluate the Awareness of Tool-augmented LLMs
To simulate incomplete scenarios where necessary tools are unavailable or users provide partial information, we manipulate instances from a test set of two benchmarks: APIBank (Li et al., 2023) and ToolBench (Qin et al., 2023). These datasets are designed to evaluate how effectively LLMs can respond to user requests using APIs. [^2] We refer to Tool and API interchangeably. In their original dataset, each instance has an available API that can address the user’s request, with sufficient information provided to invoke the API. This section describes our data source (section 3.1), manipulation strategy (section 3.2), and human verification process (section 3.3) for creating reliable incomplete scenarios.
@@ 3.1 Data Source
We construct a new dataset by leveraging two existing datasets: APIBank (Li et al., 2023) and ToolBench (Qin et al., 2023). APIBank consists of 73 APIs and 314 manually annotated multi-turn conversations. We select 450 instances from the test split, excluding those that require a tool-retrieval module. ToolBench is based on 16,000 real-world APIs across 49 categories from RapidAPI Hub. [^3] https://rapidapi.com/hub We utilize 764 test instances as filtered by Guo et al. (2024). These instances are used as the original data for manipulation.
@@ 3.2 Simulating Incomplete Scenarios with Instance Manipulation
We simulate two incomplete scenarios in which LLMs cannot properly invoke tools by manipulating the original dataset instances: (1) replacing relevant APIs with irrelevant ones, and (2) partially removing user utterances.
@@
We use a dense retriever to replace the appropriate APIs with similar but irrelevant ones. This manipulation ensures that the desired tool is unavailable, simulating a scenario where LLMs must decide not to use any of the provided tools. Specifically, the relevant APIs in the original instance are replaced with other APIs from separate API pools. To find semantically similar APIs, we use the sentence encoder proposed by Gao et al. (2021).444princeton-nlp/sup-simcse-roberta-large We concatenate the name and description of each API, convert this text into a fixed-size vector, and then select one of the most similar APIs by calculating cosine similarity between the relevant API and all available APIs in the dataset. Examples of successful and failed instances of this manipulation are in Fig. LABEL:fig:app:apibank_api_replacement_success to Fig. LABEL:fig:app:toolbench_api_replacement_wrong.
@@
We partially remove user utterances in a conversation to mimic scenarios where (1) the request is unclear, making it impossible to use APIs, or (2) the request is clear but lacks necessary information for API invocation. This manipulation removes essential information from user utterances, rendering appropriate API invocation infeasible. We automate this process using a proprietary LLM (i.e., GPT-4 (Achiam et al., 2023)) to generate naturally corrupted dialogues. The model is instructed to identify and remove critical information required for tool invocation, and we enhance the quality of this process with reasoning prompts and five manually designed few-shot samples. Sample instances are in Fig. LABEL:fig:app:apibank_utterance_removal_success to Fig. LABEL:fig:app:toolbench_utterance_removal_wrong.
Table 1: Dataset Statistics. The API Replacement and Utterance Removal denote Relevant API Replacement and Partial Removal of User Utterance, respectively. Avg. API Length denotes the string length of each API.
@@ 3.3 Data Verification
To ensure the validity of our dataset, we manually reviewed all instances to remove cases where user requests could still be handled by the provided APIs or where the manipulations resulted in unnatural conversations. Two authors holding bachelor’s degree or higher in Computer Science conducted this verification process, focusing on two key aspects: (1) whether alternative or non-replaced APIs could still fulfill the user’s request in API Replacement cases, and (2) whether the remaining information was sufficient for API execution in Utterance Removal cases. [^5] Detailed dataset verification criteria are in Appendix A.2. Through active discussion and refinement of filtering criteria, we ensured that the final dataset includes only instances that genuinely represent incomplete scenarios.
@@ 4 Experimental Setup
Phi-3 small Mistral 7B Llama3.1 8B Llama3.1 70B Qwen2 7B Claude3 haiku GPT-3.5 Turbo GPT-4 GPT-4o mini GPT-4o Relevant API Replacement APIBank Acc. 0-shot 61.60 69.38 76.06 79.10 68.77 68.89 67.92 87.85 88.82 86.03 4-shot 61.97 75.09 75.70 82.50 71.81 65.74 71.32 85.66 85.66 83.35 F1 0-shot 36.80 62.16 71.32 81.22 58.62 63.53 62.39 87.56 88.41 86.23 4-shot 38.51 69.81 74.81 83.33 63.41 47.58 65.80 84.99 85.99 82.90 ToolBench Acc. 0-shot 54.84 60.11 69.45 72.64 67.36 68.57 64.29 72.31 77.69 76.48 4-shot 55.49 60.88 76.37 80.55 68.35 58.57 57.91 79.01 75.82 77.14 F1 0-shot 14.55 34.36 56.43 76.97 52.48 59.03 46.28 61.93 72.97 75.06 4-shot 17.52 35.04 70.34 80.49 60.44 27.36 25.34 77.02 69.36 76.89 Partial Removal of User Utterance APIBank Acc. 0-shot 51.95 52.63 52.63 77.42 57.39 62.65 56.20 89.47 77.25 88.29 4-shot 62.14 52.63 59.25 84.72 57.22 53.14 65.87 88.79 79.46 89.81 F1 0-shot 4.07 23.14 19.60 80.06 29.69 52.59 36.76 89.27 73.41 88.74 4-shot 38.23 20.51 46.90 85.98 29.21 9.80 55.43 88.70 79.10 90.07 ToolBench Acc. 0-shot 49.94 50.19 50.19 71.54 51.10 53.17 49.42 57.05 57.57 74.51 4-shot 50.06 50.45 51.36 72.83 54.98 49.94 50.32 66.11 55.11 75.16 F1 0-shot 0.51 5.41 5.41 74.94 10.85 27.89 2.98 27.19 35.18 70.90 4-shot 1.03 4.49 14.93 70.25 33.59 1.02 2.04 55.74 25.05 73.33
Table 2: Performance evaluation results of LLM by manipulation type. The accuracy (Acc.) and F1 score (F1) are used for evaluation metrics. Both the zero-shot and four-shot performance are presented. The highest and the second-highest scores in each metric are highlighted in bold and underlined.
@@
We set the task where LLMs recognize incomplete conditions as a binary classification problem. We ask the LLM to determine whether APIs can be invoked to fulfill user requests based on conversations and available APIs. To assess whether LLMs can recognize both complete and incomplete conditions, we utilize the instances of two datasets(APIBank and ToolBench) and corresponding manipulated samples. LLMs should answer "Yes" or "No" for the original and manipulated instances, respectively. We measure performance using accuracy and F1 score, with "No" designated as the positive class.
@@
Following open-source and proprietary LLMs are used for our experiments: Phi-3-small-8k-instruct (Abdin et al., 2024), Mistral-Instruct-v0.2 (7B) (Jiang et al., 2023), Llama-3.1-8B/80B-Instruct (Dubey et al., 2024), Qwen2-7B-Instruct (Yang et al., 2024), Claude-3-Haiku (Anthropic, 2024), GPT-3.5-Turbo (OpenAI, 2023a), and GPT-4 (Achiam et al., 2023), GPT-4o-mini (Achiam et al., 2023), and GPT-4o (Achiam et al., 2023). [^6] More implementation details are in Appendix A.1.
5 Results and Analysis
We present our key observations as follows. Further analyses, such as the impact of tool invocation feedback, are provided in Appendix B.
Model performance varies significantly across manipulation types and datasets. Experimental results in Table 2 show distinct performance patterns. In API Replacement, GPT-4 variants and Llama 3.1 70B achieve strong results (88.82% and 79.10% accuracy on APIBank). The performance gap widens in the Utterance Removal type, especially in ToolBench, where smaller models show significant degradation. Notably, extremely low F1 scores in ToolBench stem from models’ predictions concentrating on "Yes" responses, indicating their tendency to overestimate tool applicability and suggesting that Utterance Removal requires more advanced language understanding.
Few-shot learning shows dataset-dependent effectiveness. The impact of few-shot learning varies across datasets. While models generally perform better with few-shot examples on APIBank, improvements are less pronounced in ToolBench. We attribute this to the complexity of real-world tools, which require significant reasoning skills and involve diverse scenarios (Khot et al., 2022).
Figure 2: F1-score comparison across different manipulation types and datasets, showing the impact of CoT prompting. Full results are in Table LABEL:tab:cot_full.
Chain-of-Thought prompting benefits smaller models significantly. As shown in Fig. 2, CoT prompting enhances model performance, particularly for weak models. In APIBank’s Utterance Removal task, Phi-3 small’s F1 score increases from 4.07% to 38.23% with CoT, suggesting that explicit reasoning steps facilitate a better understanding of API-instruction relationships.
Human performance establishes a reliable benchmark. We conducted a human evaluation with nine participants analyzing 20 balanced instances each from 180 selected APIBank samples. Results in Table 3 show consistent human performance (81.11% accuracy) across manipulation types, with F1 scores from 78.48% to 81.32%. This provides a meaningful reference point for model evaluation, with some larger models demonstrating comparable or superior performance in handling incomplete situations. More details on human evaluations are in Appendix A.3.